If I invested $6,000 in 1993 at 3% interest, how much do I have twenty years later?
Answer by stanbon(75887)

If I invested $6,000 in 1993 at 3% interest, how much do I have twenty years later?
-----
Assuming simple interest you get::
A(t) = P(1+rt)
-----
A(20) = 6000(1+0.03*20)
------
A(20) = 6000(1.6) = $9600
----------------------------
Cheers,
Stan H.
============================